# Qubit-Efficient QUBO Formulation for Constrained Optimization Problems

### v1

@@ 4.1 Class of Exponential Penalty Functions
The class of exponential functions that was proposed by Allran and Johnsen [17] is
$F_{k}(x)=f(x)-\sum_{i=1}^{m}\frac{1}{s(k)}e^{r(k)g_{i}(x)}$ (13)
where $f(x)$ is the objective function, $\geq s(k)\geq 1, r(k)r(k)\rightarrow\infty$ and $m$ is the number of inequality functions. From this general form, we investigate three representative penalty types:
$F_{1}=f(x)-\sum_{i=1}^{m}e^{kg_{i}(x)}$ (14)
$F_{2}=f(x)-\sum_{i=1}^{m}\frac{1}{a^{k}}e^{a^{k}g_{i}(x)}$ (15)
when $a>1$ and
$F_{3}=f(x)-\sum_{i=1}^{m}\frac{1}{a^{k}}e^{b^{k}g_{i}(x)}$ (16)
when $1<a<b$. $F_{1}$ allows steep growth for large violations, suitable when strict constraint enforcement is required. $F_{2}$ controls both magnitude and steepness, balancing early and late penalty effects. And $F_{3}$ allows independent tuning of initial and late-stage penalty growth, offering maximum flexibility. The parameters $a,b,k$ control how quickly the penalty ramps up, enabling adaptation to different constraint tightness.
Interior penalty methods penalize violations within the feasible region, exterior penalty methods penalize the infeasible solutions. A notable disadvantage of interior penalty functions is the potential difficulty in finding an interior starting solution, as this process can require as much effort as solving the entire nonlinear program itself. Additionally, selecting appropriate penalty parameters is critical, as improper choices can lead to slow convergence or numerical instability, especially near the optimal solution. These challenges are further compounded in large-scale or ill-conditioned problems, where stability and efficiency become critical concerns.
Another issue arises in handling non-convex problems, where the methods’ sensitivity to initial solutions can result in convergence to local optima rather than the global optimum. While trial solutions for exponential penalty functions can be feasible or infeasible, ensuring feasibility through parameter selection may still be problematic in cases of highly restrictive or complex constraints. Furthermore, interior penalty methods may suffer from inefficiency when compared to alternative approaches, such as barrier methods or exterior penalty functions, which are often better suited to certain problem structures.
Despite these limitations, interior penalty functions remain useful in many convex cases, where appropriate parameter tuning ensures the feasibility of trial solutions and convergence to the optimal solution. However, their application in more general nonlinear programming problems requires careful consideration of these trade-offs.
@@
The class of exponential penalty functions has the property that they require uniformly bounded over any compact feasible region, and again, by an appropriate choice of parameters, the gradient remains uniformly bounded over the feasible region. Here as $r(k)\rightarrow\infty$, the function $F_{k}(x)$ in 13 has the value $f(x)$ if $x$ well satisfies all the constraints. If $s(k)\rightarrow\infty$, this is true even for the points at the boundary of the feasible region. As with other penalty functions, it can be a concave function, ensuring that a local maximum is a global maximum.
@@
By comparing exponential and slack approaches empirically, we derived the following statements that show the relation between the number of variables in problem and required qubits:
@@ 5 Experimental Setup
This section details the experimental framework employed to assess the performance of the proposed exponential penalty encoding method. We describe the experimental setup, evaluation metrics, hyperparameter selection, and dataset selection.
@@ 5.1 Evaluation Metrics
We evaluate how the number of required qubits $Q(n)$ grows with problem size $n$. To quantify improvements over baseline approaches, we define the qubit reduction ratio as:
@@ 5.3 Dataset and Experimental Testbed
For the implementation and analysis, we use Python 3.11 and Qiskit on Lenovo ThinkPad X1 Yoga Intel(R) Core(TM) i7-8550U CPU @ 1.80GHz. The experiments were performed on server AMD EPYC 7452 32-Core Processor with 128 CPUs, 2 sockets, 32 cores per socket, and 2 threads per core.
@@ 6.3 Approximation Probability
Figure 4: The mean convergence of QAOA with exponential encoding for single problem samples executed on noiseless simulators.
@@ 6.4 Convergence Time
Convergence time for BPP with 8 qubits and TSP with 12 qubits instances are on average 7 and 23 seconds respectively. While the slack variable approach requires a considerably higher amount of qubits, we do not consider the slack variable approach in convergence time analysis. QAOA execution of large-scale samples that exceed the number of qubits 20, takes too long to run. For example, BPP sample with bins 2, items 3 requires 8 qubits with exponential encoding, while with slack encoding, it requires 24 qubits. Relatively, simulating QAOA solution of such instance on noise-free quantum computer takes around 5 seconds with exponential encoding and over 6 hours with slack encoding as it requires 24 qubits.

### v2

@@ 4.1 Class of Exponential Penalty Functions
We define a general exponential penalty formulation
$F_{k}(x)=f(x)-\sum_{i=1}^{m}P_{k}\big(g_{i}(x)\big),$ (13)
where $P_{k}:\mathbb{R}\to\mathbb{R}$ is a penalty term parameterized by $k$ and $f(x)$ is the objective function corresponding to objective funtion in Equation 1. We study three exponential forms with distinct scaling properties. For $k\in\mathbb{N}$ and constants $a>1$ and $b>a$, define:
$P_{k}^{(1)}(g_{i}(x))$ $=e^{kg_{i}(x)},$ (14)
$P_{k}^{(2)}(g_{i}(x))$ $=\frac{1}{a^{k}}\,e^{a^{k}g_{i}(x)},$ (15)
$P_{k}^{(3)}(g_{i}(x))$ $=\frac{1}{a^{k}}\,e^{b^{k}g_{i}(x)}.$ (16)
where $g_{i}(x)$ is an inequality term 3. The corresponding penalized objectives are $F_{r}(x)=f(x)-\sum_{i=1}^{m}P_{k}^{(r)}(g_{i}(x))$, for $r\in\{1,2,3\}$.
4.2 Simple exponential penalty $P^{1}$
The penalty in 14 introduces a direct exponential encoding of constraint violation and is fully governed by scalar parameter $k>0$:
$F_{1}(x)=f(x)-\sum_{i=1}^{m}e^{kg_{i}(x)},$ (17)
The exponential term $e^{kg_{i}(x)}$ exhibits the following properties:
  1. 1.
    Curvature: The first and second derivatives with respect to $g_{i}(x)$ are
    $\frac{d}{dg_{i}}e^{kg_{i}(x)}=ke^{kg_{i}(x)},\quad\frac{d^{2}}{dg_{i}^{2}}e^{kg_{i}(x)}=k^{2}e^{kg_{i}(x)}.$
    Both are strictly positive, implying that the penalty is smooth, convex, and strictly increasing in $g_{i}(x)$.
  2. 2.
    Boundary behavior: For $g_{i}(x)\leq 0$, the penalty remains bounded above by 1. Near the constraint boundary, a second-order Taylor expansion gives
    $e^{kg_{i}(x)}=1+kg_{i}(x)+\frac{1}{2}k^{2}g_{i}(x)^{2}+O(g_{i}(x)^{3}),$
    which shows that local curvature increases quadratically with $k$. Thus, large $k$ values result in rapid growth even for small violations, making this form highly sensitive near the feasible boundary.
    For any fixed violation $g_{i}(x)>0$, $e^{kg_{i}(x)}\to\infty$ as $k\to\infty$. Consequently, infeasible solutions become unbounded in the objective, ensuring convergence toward the feasible region.
4.3 Scaled exponential penalty $P^{2}$
The penalty in 15 introduces a scale normalization to balance the early- and late-stage effects of the exponential term:
$F_{2}(x)=f(x)-\sum_{i=1}^{m}\frac{1}{a^{k}}e^{a^{k}g_{i}(x)},\quad a>1.$ (18)
Here, the term $a^{-k}$ counteracts part of the exponential growth, providing a softer penalty when $g_{i}(x)$ is close to zero, while still ensuring divergence for large violations.
  1. 1.
    Boundary behavior: Expanding around $g_{i}(x)=0$,
    $\frac{1}{a^{k}}e^{a^{k}g_{i}(x)}=\frac{1}{a^{k}}\left(1+a^{k}g_{i}(x)+\frac{(a^{k}g_{i}(x))^{2}}{2}+O(g_{i}(x)^{3})\right)=\frac{1}{a^{k}}+g_{i}(x)+\frac{a^{k}g_{i}(x)^{2}}{2}+O(g_{i}(x)^{3}).$
    The linear term is independent of $k$, implying that near the feasible region the penalty responds linearly to small violations, regardless of the penalty strength. This allows gentle exploration around the constraint boundary without excessive gradient magnitudes.
    For $g_{i}(x)>0$, the exponential term dominates:
    $\frac{1}{a^{k}}e^{a^{k}g_{i}(x)}\approx a^{-k}e^{a^{k}g_{i}(x)}\to\infty\quad\text{as }k\to\infty,$
    so feasibility is still asymptotically enforced.
  2. 2.
    Curvature: The gradient contribution from the penalty term is $e^{a^{k}g_{i}(x)}\nabla g_{i}(x)$, which remains modest near feasibility and increases sharply when constraints are violated. The corresponding curvature term scales as $a^{k}e^{a^{k}g_{i}(x)}$, making the landscape steeper only for substantial violations.
4.4 Dual-rate exponential panalty $P^{3}$
The penalty $P^{3}$ introduces two scaling parameters, $a$ and $b$, for early-stage smoothness and late-stage enforcement:
$F_{3}(x)=f(x)-\sum_{i=1}^{m}\frac{1}{a^{k}}e^{b^{k}g_{i}(x)},\quad 1<a<b.$ (19)
This construction generalizes $F_{2}$ by allowing independent control over the penalty’s growth rate inside and outside the feasible region.
  1. 1.
    Boundary behavior: Near $g_{i}(x)=0$,
    $\frac{1}{a^{k}}e^{b^{k}g_{i}(x)}=\frac{1}{a^{k}}\left(1+b^{k}g_{i}(x)+\frac{1}{2}b^{2k}g_{i}(x)^{2}+O(g_{i}(x)^{3})\right)=\frac{1}{a^{k}}+\frac{b^{k}}{a^{k}}g_{i}(x)+\frac{b^{2k}}{2a^{k}}g_{i}(x)^{2}+O(g_{i}(x)^{3}).$
    The linear term grows as $(b/a)^{k}$, so increasing $b/a$ intensifies the gradient pressure near feasibility, allowing flexible adjustment of local sensitivity.
    For $g_{i}(x)>0$,
    $\frac{1}{a^{k}}e^{b^{k}g_{i}(x)}\to\infty,$
    at a rate exponentially faster than $F_{2}$, since $b^{k}>a^{k}$. Hence, constraint violations are penalized more aggressively, ensuring faster convergence toward feasibility in later stages.
  2. 2.
    Curvature: The gradient contribution is $(b^{k}/a^{k})e^{b^{k}g_{i}(x)}\nabla g_{i}(x)$, and the curvature scales as $(b^{2k}/a^{k})e^{b^{k}g_{i}(x)}$. This formulation allows one to control both the magnitude and sharpness of the penalty through the ratio $b/a$.
@@ 4.4 Dual-rate exponential panalty $P^{3}$
These exponential penalty functions has the property that they are uniformly bounded over any compact feasible region, and again, by an appropriate choice of parameters, the gradient remains uniformly bounded over the feasible region.
@@
By comparing the general form of exponential function and slack approaches empirically, we derived the following statements that show the relation between the number of variables in the problem and required qubits:
@@ 5 Experimental Setup
This section details the experimental framework employed to assess the performance of the proposed exponential penalty encoding method. We describe the experimental setup, evaluation metrics, hyperparameters, and dataset selection.
@@ 5.1 Evaluation Metrics
We evaluate how the number of required qubits $Q(n)$ grows with problem size $n$. To quantify improvements over baseline approach $Q_{\text{slack}}$, we define the qubit reduction ratio as:
@@ 5.3 Dataset and Experimental Testbed
For the implementation and analysis, we use Python 3.11 and Qiskit on a Lenovo ThinkPad X1 Yoga Intel(R) Core(TM) i7-8550U CPU @ 1.80GHz. The experiments were performed on server AMD EPYC 7452 32-Core Processor with 128 CPUs, 2 sockets, 32 cores per socket, and 2 threads per core.
@@ 6.3 Approximation Probability
Figure 4: The mean convergence of QAOA with exponential encoding for single problem samples executed on noiseless simulators with hyperparameters $r=10,s=5,p=2$.
@@ 6.4 Convergence Time
Convergence time for BPP with 8 qubits and TSP with 12 qubits instances are, on average, 7 and 23 seconds, respectively. While the slack variable approach requires a considerably higher amount of qubits, we do not consider the slack variable approach in the convergence time analysis. QAOA execution of large-scale samples that exceed 20 qubits takes too long to run. For example, BPP sample with bins 2, items 3 requires 8 qubits with exponential encoding, while with slack encoding, it requires 24 qubits. Relatively, simulating the QAOA solution of such an instance on a noise-free quantum computer takes around 5 seconds with exponential encoding and over 6 hours with slack encoding, as it requires 24 qubits.